# Leaf schemes and Hodge loci [^1] February 25, 2025

### v1

@@ 4.5 Final remarks
As the main goal of the present article has been to prepare the ground for similar investigations in the case of Calabi-Yau varieties, one might try to describe modulo prime properties of topological string partition functions $F_{g}$ in the case of elliptic curves, see the articles of Dijkgraaf, Douglas, Kaneko and Zagier in [Mov12, Appendix B]. These are homogeneous polynomials in $E_{2},E_{4}$ and $E_{6}$ of weight $66g-$ 6, for instance $F_{2}=\frac{1}{103680}(10E_{2}^{3}-6E_{2}E_{4}-4E_{6})$. For Calabi-Yau threefolds, and in particular for mirror quintic, there is an ambiguity problem for $F_{g}$’s which has been only established for lower genus, see the references in [Mov17a]. Modulo primes investigation of $F_{g}$’s might give some insight to this problem. For instance, in the last lines of the computer code of 4.2 we have investigated the zero locus of $F_{2}$ restricted to the curve ${\rm Zero}({{\cal I}}_{p})$. We found many such points with coordinates in $\mathbb{F}_{p}$, but no pattern for different primes were found.
@@ 5.1 Introduction
Let ${\sf T}$ be the space of homogeneous polynomials $f(x)$ of degree $d$ in $n+2$ variables $x=(x_{0},x_{1},\ldots,x_{n+1})$ and with coefficients in $\mathbb{C}$ such that the induced hypersurface $X:=\mathbb{P}\{f=0\}$ in $\mathbb{P}^{n+1}$ is smooth. We assume that $n\geq 2$ is even and $d\geq 3$. Consider the subvariety of ${\sf T}$ parametrizing hypersurfaces containing two projective subspaces $\mathbb{P}^{\frac{n}{2}},\check{\mathbb{P}}^{\frac{n}{2}}$ (we call them linear cycles) with $\mathbb{P}^{\frac{n}{2}}\cap\check{\mathbb{P}}^{\frac{n}{2}}=\mathbb{P}^{m}$ for a fixed $-1\leq m\leq\frac{n}{2}-1$ ($\mathbb{P}^{-1}$ is the empty set). We are actually interested in a local analytic branch $V_{Z}$ of this space which parametrizes deformations of a fixed $X$ together with such two linear cycles. We consider the algebraic cycle
(54) $Z=r\mathbb{P}^{\frac{n}{2}}+\check{r}\check{\mathbb{P}}^{\frac{n}{2}},\ \ r\in\mathbb{N},0\not=\check{r}\in\mathbb{Z}$
@@
Note that $V_{Z}$ does not depend on $r$ and $\check{r}$ and it is $V_{Z}=V_{\mathbb{P}^{\frac{n}{2}}}\cap V_{\check{\mathbb{P}}^{\frac{n}{2}}}$, where $V_{\mathbb{P}^{\frac{n}{2}}}$ and $V_{\check{\mathbb{P}}^{\frac{n}{2}}}$ are two branches of the subvariety of ${\sf T}$ parameterizing hypersurfaces containing a linear cycle, see Figure 2. From now on we use the notation $t\in{\sf T}$ and denote the corresponding polynomial and hypersurface by $f_{t}$ and $X_{t}$ respectively, being clear that $f_{0}=f$ and $X_{0}=X$. The monodromy/parallel transport $\delta_{t}\in H^{n}(X_{t},\mathbb{Z})$ is well-defined for all $t\in({\sf T},0)$, a small neighborhood of $t$ in ${\sf T}$ with the usual/analytic topology, and it is not necessarily supported in algebraic cycles like the original $\delta_{0}$. We arrive at the set theoretical definition of the Hodge locus
(55) $V_{[Z]}:=\left\{t\in({\sf T},0)\ \ |\ \ \delta_{t}\hbox{ is a Hodge cycle, that is }\ \ \delta_{t}\in H^{\frac{n}{2},\frac{n}{2}}(X_{t})\cap H^{n}(X_{t},\mathbb{Z})\right\}.$
@@
For $n=4$ the Hodge conjecture is a theorem and the first part of the above conjecture for $n=4$ is true for trivial reasons. If the first part of the above conjecture is true then one might try to verify the Hodge conjecture for the Hodge cycle $\delta_{t},\ t\in V_{[Z]}$ which is absolute, see Deligne’s lecture in [DMOS82]. It is only verified for $t\in V_{Z}$ using the algebraic cycle $Z$. By Cattani-Deligne-Kaplan theorem $V_{[Z]}$ for fixed $r$ and $\check{r}$ is a union of branches of an algebraic set in ${\sf T}$ and we will have the challenge of verifying a particular case of Grothendieck’s variational Hodge conjecture. It can be verified easily that the tangent spaces of $V_{[Z]}$ intersect each other in the tangent space of $V_{Z}$, and hence, we get a pencil of Hodge loci depending on the rational number $\frac{r}{\check{r}}$, see Figure 2. Similar computations as for 5.1 in the case of surfaces result in a conjectural counterexample to a conjecture of J. Harris for degree $8$ surfaces, see [Mov21b].
@@
Figure 2: A pencil of Hodge loci
@@
The computational methods introduced in [Mov21a] can be applied to an arbitrary combination of linear cycles, for some examples see [Movxx, Chapter 1], however, for simplicity the author focused mainly in the sum of two linear cycles as announced earlier. We note that $V_{[Z]}$ carries a natural analytic scheme/space structure, that is, there is an ideal $I=\langle f_{1},f_{2},\ldots,f_{k}\rangle\subset{\cal O}_{{\sf T},0}$ of holomorphic functions $f_{i}$ in a small neighborhood $({\sf T},0)$ of 00 in ${\sf T}$, and the ring structure of $V_{[Z]}$ is ${\cal O}_{{\sf T},0}/I$. The holomorphic functions $f_{i}$ are periods $\int_{\delta_{t}}\omega_{i}$, where $\omega_{i}$’s are global sections of the $n$-th cohomology bundle $\cup_{t\in({\sf T},0)}H^{n}_{\rm dR}(X_{t})$ such that for fixed $t$ they form a basis of the piece $F^{\frac{n}{2}+1}H^{n}_{\rm dR}(X_{t})$ of Hodge filtration (form now on all Hodge cycles will be considered in homology and not cohomology). For hypersurfaces, using Griffiths work [Gri69], the holomorphic functions $f_{i}$’s are
(56) ${{\int}}_{\delta_{t}}{\rm Resi}\left(\frac{x^{\beta}\Omega}{f^{k}_{t}}\right),\ \ k=1,2,\cdots,\frac{n}{2},\ \ \ x^{\beta}\in(\mathbb{C}[x]/{\rm jacob}(f_{t}))_{kd-n-2}$
and $x^{\beta}$ is a basis of monomials for the degree $kd-n-2$ piece of the Jacobian ring $\mathbb{C}[x]/{\rm jacob}(f_{t})$ and $\Omega:=\sum_{i=0}^{n+1}x_{i}dx_{0}\wedge dx_{1}\wedge\cdots\wedge dx_{i-1}\wedge dx_{i+1}\wedge\cdots\wedge dx_{n+1}$. The Taylor series of such integrals can be computed and implemented in a computer, however, for simplicity we have done this around the Fermat variety.
@@
(57) $f_{t}:=x_{0}^{d}+x_{1}^{d}+\cdots+x_{n+1}^{d}-\sum_{\alpha}t_{\alpha}x^{\alpha}=0,$
@@
where $\alpha$ runs through a finite subset $I$ of $\mathbb{N}_{0}^{n+2}$ with $\sum_{i=0}^{n+1}\alpha_{i}=d$. From now on for all statements and conjectures $X_{0}$ is the Fermat variety. The Taylor series for the Fermat variety $X_{0}$ can be computed explicitly, see [Mov21a, 18.5].  [^18] For particular cases of such series see (16) and (27). It is also implemented in computer, see [Mov21a, Section 20.11]. Its announcement takes almost a full page and we only content ourselves to the following statement:
@@
Let $\delta_{0}\in H_{n}(X_{0},\mathbb{Q})$ be a Hodge cycle and $x^{\beta}$ be a monomial of degree $kd-n-2$. The integral $\frac{1}{(2\pi i)^{\frac{n}{2}}}{{\int}}_{\delta_{t}}{\rm Resi}\left(\frac{x^{\beta}\Omega}{f^{k}_{t}}\right)$ can be written as a power series in $(t_{\alpha})_{\alpha\in I}$ with coefficients in an abelian extension of $\mathbb{Q}(\zeta_{d})$. If $\delta_{0}$ is a sum of linear cycles $\mathbb{P}^{\frac{n}{2}}$ then such an abelian extension is $\mathbb{Q}(\zeta_{2d})$.
In 5.1 we have considered $V_{[Z]}$ as an analytic variety. As an analytic scheme and for $X_{0}$ the Fermat variety, we even claim that $V_{[Z]}$ is smooth which implies that it is also reduced. The first goal is to compare the dimension of Zariski tangent spaces ${\rm dim}(T_{t}V_{[Z]})$ and ${\rm dim}(T_{t}V_{Z})$. Computation of $TV_{[Z]}$ is done using the notion of infinitesimal variation of Hodge structures developed by P. Griffiths and his coauthors in [CGGH83]. In a down-to-earth terms, this is just the data of the linear parts of $f_{i}$’s. It turns out that
@@
This is proved in [Mov21a, Theorem 18.1] for
(58) $0<r\leq|\check{r}|\leq 10$
@@
$(2,d),\ \ \ d\leq 14,\ \ (4,3),(4,4),(4,5),(4,6),\ (6,3),(6,4),(8,3),(8,3),(10,3),(10,3),(10,3),$
using computer. For the proof of 5.1 we have computed both ${\rm dim}T_{0}V_{[Z]}$ and ${\rm dim}(V_{Z})$ and we have verified that these dimensions are equal. The full proof of 5.1 is done in [VL22b, Theorem 1.3]. Throughout the paper, the condition (58) is needed for all statements whose proof uses computer, however, note that the number $10$ is just the limit of the computer and the author’s patience for waiting the computer produces results. All the conjectures that will appear in this section are not considered to be so difficult and their proofs or disproofs are in the range of available methods in the literature.
@@
In [Mov21a, Theorem 18.3, part 1] we have proved the above conjecture by computer for $(n,d)$ in the list
[…]
see also [Dan17b] for many examples of this situation in the case of surfaces, that is, $n=2$.
@@
Note that in the situation of 5.2, $\mathbb{P}^{\frac{n}{2}}+\check{\mathbb{P}}^{\frac{n}{2}}$ is a complete intersection of the mentioned type. In this way 5.2 follows from [Dan17a], see also [MV21, Chapter 11]. In our search for a Hodge locus $V_{[Z]}$ bigger than $V_{Z}$ we arrive at the cases
@@
This conjecture for $n=6,8$ is proved in [Mov21a, Theorem 18.3 part 2]. The same conjecture for $(n,d,m)=(4,4,0)$ is also proved there.
@@
For $(d,m)=(3,\frac{n}{2}-2)$ with $(r,\check{r})=(1,-1)$, the Hodge locus $V_{[Z]}$ is smooth and it parameterizes hypersurfaces containing generalized cubic scroll (for the definition see Section 5.5 and [Mov21a, Section 19.6]) .
This conjecture is obtained after a series of email discussions with P. Deligne in 2018, see [Movxx, Chapter 1] and [Mov21a, Section 19.6]. The proof of this must not be difficult (comparing two tangent spaces). The case $(n,d,m)=(4,4,0)$ with $(r,\check{r})=(1,-1)$ is still mysterious, however, it might be solved by similar methods as in the mentioned references. The only remaining cases are the case of 5.1 for arbitrary even number $n\geq 4$ and $(d,m)=(4,\frac{n}{2}-2),\ n\geq 6$. It turns out 5.1 is false for $n\geq 10$, see Section 5.9.
@@ 5.3 Evidence 1
The first evidence to 5.1 comes from computing the Zariski tangent spaces of both $V_{Z}$ and $V_{[Z]}$, for the Fermat variety $X_{0}$, and observing that ${\rm dim}({\sf T}_{t}V_{[Z]})={\rm dim}({\sf T}_{t}V_{Z})+1$. This has been verified by computer for many examples of $n$ in [Mov21a, Chapter 19] and the full proof can be found in Section 5.10. However, this is not sufficient as $V_{[Z]}$ carries a natural analytic scheme structure. Moreover, $V_{[Z]}$ as a variety might be singular, even though, the author is not aware of an example. The Zariski tangent space is only the first approximation of a variety, and one can introduce the $N$-th order approximations $V^{N}_{[Z]},\ \ N\geq 1$ which we call it the $N$-th infinitesimal Hodge locus, such that $V^{1}_{[Z]}$ is the Zariski tangent space. The algebraic variety $V^{N}_{[Z]}$ is obtained by truncating the defining holomorphic functions of $V_{Z}$ up to degree $N$. The non-smoothness results as above follows from the non-smoothness of $V^{N}_{[Z]}$ for small values of $N$ like $2,3$ (the case $N=2$ has been partially treated in cohomological terms in [Mac05]). The strongest evidence to 5.1 is the following theorem in [Mov21a, Theorem 19.1, part 2] which is proved by heavy computer calculations.
@@
[…]
For $n=4$, the Hodge locus $V_{[Z]}$ itself is smooth for trivial reasons. There is abundant examples of Hodge cycles for which we know neither to verify the Hodge conjecture (construct algebraic cycles) nor give evidences that they might be counterexamples to the Hodge conjecture, see [Del06] and [Mov21a, Chapter 19]. Finding Hodge cycles for hypersurfaces is extremely difficult, and the main examples in this case are due to T. Shioda for Fermat varieties [Shi79].

### v2

@@ 4.7 Final remarks
As the main goal of the present article has been to prepare the ground for similar investigations in the case of Calabi-Yau varieties, one might try to describe modulo prime properties of topological string partition functions $F_{g}$ in the case of elliptic curves, see the articles of Dijkgraaf, Douglas, Kaneko and Zagier in [ho14, Appendix B]. These are homogeneous polynomials in $E_{2},E_{4}$ and $E_{6}$ of weight 6g66g-6, for instance $F_{2}=\frac{1}{103680}(10E_{2}^{3}-6E_{2}E_{4}-4E_{6})$. For Calabi-Yau threefolds, and in particular for mirror quintic, there is an ambiguity problem for $F_{g}$’s which has been only established for lower genus, see the references in [GMCD-MQCY3]. Modulo primes investigation of $F_{g}$’s might give some insight to this problem. For instance, in the last lines of the computer code of 4.2 we have investigated the zero locus of $F_{2}$ restricted to the curve ${\rm Zero}({{\mathcal{I}}}_{p})$. We found many such points with coordinates in $\mathbb{F}_{p}$, but no pattern for different primes were found.
@@ 5.1 Introduction
Let $\sf T$} be the space of homogeneous polynomials $f(x)$ of degree $d$ in $n+2$ variables $x=(x_{0},x_{1},\ldots,x_{n+1})$ and with coefficients in $\mathbb{C}$ such that the induced hypersurface $X:=\mathbb{P}\{f=0\}$ in $\mathbb{P}^{n+1}$ is smooth. We assume that $n\geq 2$ is even and $d\geq 3$. Consider the subvariety of $\sf T$} parametrizing hypersurfaces containing two projective subspaces $\mathbb{P}^{\frac{n}{2}},\check{\mathbb{P}}^{\frac{n}{2}}$ (we call them linear cycles) with $\mathbb{P}^{\frac{n}{2}}\cap\check{\mathbb{P}}^{\frac{n}{2}}=\mathbb{P}^{m}$ for a fixed $1\leq m\leq\frac{n}{2}-1$ ($\mathbb{P}^{-1}$ is the empty set). We are actually interested in a local analytic branch $V_{Z}$ of this space which parametrizes deformations of a fixed $X$ together with such two linear cycles. We consider the algebraic cycle
(5.1) $Z=r\mathbb{P}^{\frac{n}{2}}+\check{r}\check{\mathbb{P}}^{\frac{n}{2}},\ \ r\in\mathbb{N},0\not=\check{r}\in\mathbb{Z}$
@@
Note that $V_{Z}$ does not depend on $r$ and $\check{r}$ and it is $V_{Z}=V_{\mathbb{P}^{\frac{n}{2}}}\cap V_{\check{\mathbb{P}}^{\frac{n}{2}}}$, where $V_{\mathbb{P}^{\frac{n}{2}}}$ and $V_{\check{\mathbb{P}}^{\frac{n}{2}}}$ are two branches of the subvariety of $\sf T$} parameterizing hypersurfaces containing a linear cycle, see Figure 5.1. From now on we use the notation $t\in{\sf T}$ and denote the corresponding polynomial and hypersurface by $f_{t}$ and $X_{t}$ respectively, being clear that $f_{0}=f$ and $X_{0}=X$. The monodromy/parallel transport $\delta_{t}\in H^{n}(X_{t},\mathbb{Z})$ is well-defined for all $t\in({\sf T},0)$, a small neighborhood of $t$ in $\sf T$} with the usual/analytic topology, and it is not necessarily supported in algebraic cycles like the original $\delta_{0}$. We arrive at the set theoretical definition of the Hodge locus
(5.2) $V_{[Z]}:=\left\{t\in({\sf T},0)\ \ |\ \ \delta_{t}\hbox{ is a Hodge cycle, that is }\ \ \delta_{t}\in H^{\frac{n}{2},\frac{n}{2}}(X_{t})\cap H^{n}(X_{t},\mathbb{Z})\right\}.$
@@
For $n=4$ the Hodge conjecture is a theorem and the first part of the above conjecture for $n=4$ is true for trivial reasons. If the first part of the above conjecture is true then one might try to verify the Hodge conjecture for the Hodge cycle $\delta_{t},\ t\in V_{[Z]}$ which is absolute, see Deligne’s lecture in [dmos]. It is only verified for $t\in V_{Z}$ using the algebraic cycle $Z$. By Cattani-Deligne-Kaplan theorem $V_{[Z]}$ for fixed $r$ and $\check{r}$ is a union of branches of an algebraic set in $\sf T$} and we will have the challenge of verifying a particular case of Grothendieck’s variational Hodge conjecture. It can be verified easily that the tangent spaces of $V_{[Z]}$ intersect each other in the tangent space of $V_{Z}$, and hence, we get a pencil of Hodge loci depending on the rational number $\frac{r}{\check{r}}$, see Figure 5.1. Similar computations as for 5.1 in the case of surfaces result in a conjectural counterexample to a conjecture of J. Harris for degree $8$ surfaces, see [ho2019].
@@
Figure 5.1: A pencil of Hodge loci
@@
The computational methods introduced in [ho13] can be applied to an arbitrary combination of linear cycles, for some examples see [Headaches, Chapter 1], however, for simplicity the author focused mainly in the sum of two linear cycles as announced earlier. We note that $V_{[Z]}$ carries a natural analytic scheme/space structure, that is, there is an ideal $I=\langle f_{1},f_{2},\ldots,f_{k}\rangle\subset{\mathcal{O}}_{{\sf T},0}$ of holomorphic functions $f_{i}$ in a small neighborhood $({\sf T},0)$ of 0 in $\sf T$}, and the ring structure of $V_{[Z]}$ is ${\mathcal{O}}_{{\sf T},0}/I$. The holomorphic functions $f_{i}$ are periods $\int_{\delta_{t}}\omega_{i}$, where $\omega_{i}$’s are global sections of the $n$-th cohomology bundle $\cup_{t\in({\sf T},0)}H^{n}_{\rm dR}(X_{t})$ such that for fixed $t$ they form a basis of the piece $F^{\frac{n}{2}+1}H^{n}_{\rm dR}(X_{t})$ of Hodge filtration (form now on all Hodge cycles will be considered in homology and not cohomology). For hypersurfaces, using Griffiths work [gr69], the holomorphic functions $f_{i}$’s are
(5.3) ${{\int}}_{\delta_{t}}{\rm Resi}\left(\frac{x^{\beta}\Omega}{f^{k}_{t}}\right),\ \ k=1,2,\cdots,\frac{n}{2},\ \ \ x^{\beta}\in(\mathbb{C}[x]/{\rm jacob}(f_{t}))_{kd-n-2}$
and $x^{\beta}$ is a basis of monomials for the degree kdn2kd-n-2 piece of the Jacobian ring $\mathbb{C}[x]/{\rm jacob}(f_{t})$ and $\Omega:=\sum_{i=0}^{n+1}x_{i}dx_{0}\wedge dx_{1}\wedge\cdots\wedge dx_{i-1}\wedge dx_{i+1}\wedge\cdots\wedge dx_{n+1}$. The Taylor series of such integrals can be computed and implemented in a computer, however, for simplicity we have done this around the Fermat variety.
@@
(5.4) $f_{t}:=x_{0}^{d}+x_{1}^{d}+\cdots+x_{n+1}^{d}-\sum_{\alpha}t_{\alpha}x^{\alpha}=0,\ \$
@@
where $\alpha$ runs through a finite subset $I$ of $\mathbb{N}_{0}^{n+2}$ with $\sum_{i=0}^{n+1}\alpha_{i}=d$. From now on for all statements and conjectures $X_{0}$ is the Fermat variety. The Taylor series for the Fermat variety $X_{0}$ can be computed explicitly, see [ho13, 18.5].  [^1] For particular cases of such series see (1.16) and (3.1). It is also implemented in computer, see [ho13, Section 20.11]. Its announcement takes almost a full page and we only content ourselves to the following statement:
@@
Let $\delta_{0}\in H_{n}(X_{0},\mathbb{Q})$ be a Hodge cycle and $x^{\beta}$ be a monomial of degree kdn2kd-n-2. The integral $\frac{1}{(2\pi i)^{\frac{n}{2}}}{{\int}}_{\delta_{t}}{\rm Resi}\left(\frac{x^{\beta}\Omega}{f^{k}_{t}}\right)$ can be written as a power series in $(t_{\alpha})_{\alpha\in I}$ with coefficients in an abelian extension of $\mathbb{Q}(\zeta_{d})$. If $\delta_{0}$ is a sum of linear cycles $\mathbb{P}^{\frac{n}{2}}$ then such an abelian extension is $\mathbb{Q}(\zeta_{2d})$.
In 5.1 we have considered $V_{[Z]}$ as an analytic variety. As an analytic scheme and for $X_{0}$ the Fermat variety, we even claim that $V_{[Z]}$ is smooth which implies that it is also reduced. The first goal is to compare the dimension of Zariski tangent spaces ${\rm dim}(T_{t}V_{[Z]})$ and ${\rm dim}(T_{t}V_{Z})$. Computation of $TV_{[Z]}$ is done using the notion of infinitesimal variation of Hodge structures developed by P. Griffiths and his coauthors in [CGGH1983]. In a down-to-earth terms, this is just the data of the linear parts of $f_{i}$’s. It turns out that
@@
This is proved in [ho13, Theorem 18.1] for
(5.5) $0<r\leq|\check{r}|\leq 10$
@@
$(2,d),\ \ \ d\leq 14,\ \$$(4,3),(4,4),(4,5),(4,6),\ (6,3),(6,4),(8,3),$
$(8,3),(10,3),(10,3),(10,3),$
using computer. For the proof of 5.1 we have computed both ${\rm dim}T_{0}V_{[Z]}$ and ${\rm dim}(V_{Z})$ and we have verified that these dimensions are equal. The full proof of 5.1 is done in [RobertoThesis, Theorem 1.3]. Throughout the paper, the condition (5.5) is needed for all statements whose proof uses computer, however, note that the number $10$ is just the limit of the computer and the author’s patience for waiting the computer produces results. All the conjectures that will appear in this section are not considered to be so difficult and their proofs or disproofs are in the range of available methods in the literature.
@@
In [ho13, Theorem 18.3, part 1] we have proved the above conjecture by computer for $(n,d)$ in the list
$(2,d),\ 5\leq d\leq 9,(4,4),(4,5),(6,3),(8,3),$
see also [Dan2017] for many examples of this situation in the case of surfaces, that is, $n=2$.
@@
Note that in the situation of 5.2, $\mathbb{P}^{\frac{n}{2}}+\check{\mathbb{P}}^{\frac{n}{2}}$ is a complete intersection of the mentioned type. In this way 5.2 follows from [Dan-2014], see also [ho13-Roberto, Chapter 11]. In our search for a Hodge locus $V_{[Z]}$ bigger than $V_{Z}$ we arrive at the cases
@@
This conjecture for $n=6,8$ is proved in [ho13, Theorem 18.3 part 2]. The same conjecture for $(n,d,m)=(4,4,0)$ is also proved there.
@@
For $(d,m)=(3,\frac{n}{2}-2)$ with $(r,\check{r})=(1,-1)$, the Hodge locus $V_{[Z]}$ is smooth and it parameterizes hypersurfaces containing generalized cubic scroll (for the definition see Section 5.5 and [ho13, Section 19.6]) .
This conjecture is obtained after a series of email discussions with P. Deligne in 2018, see [Headaches, Chapter 1] and [ho13, Section 19.6]. The proof of this must not be difficult (comparing two tangent spaces). The case $(n,d,m)=(4,4,0)$ with $(r,\check{r})=(1,-1)$ is still mysterious, however, it might be solved by similar methods as in the mentioned references. The only remaining cases are the case of 5.1 for arbitrary even number $n\geq 4$ and $(d,m)=(4,\frac{n}{2}-2),\ n\geq 6$. It turns out 5.1 is false for $n\geq 10$, see Section 5.9.
@@ 5.3 Evidence 1
The first evidence to 5.1 comes from computing the Zariski tangent spaces of both $V_{Z}$ and $V_{[Z]}$, for the Fermat variety $X_{0}$, and observing that ${\rm dim}({\sf T}_{t}V_{[Z]})={\rm dim}({\sf T}_{t}V_{Z})+1$. This has been verified by computer for many examples of $n$ in [ho13, Chapter 19] and the full proof can be found in Section 5.10. However, this is not sufficient as $V_{[Z]}$ carries a natural analytic scheme structure. Moreover, $V_{[Z]}$ as a variety might be singular, even though, the author is not aware of an example. The Zariski tangent space is only the first approximation of a variety, and one can introduce the $N$-th order approximations $V^{N}_{[Z]},\ \ N\geq 1$ which we call it the $N$-th infinitesimal Hodge locus, such that $V^{1}_{[Z]}$ is the Zariski tangent space. The algebraic variety $V^{N}_{[Z]}$ is obtained by truncating the defining holomorphic functions of $V_{Z}$ up to degree $N$. The non-smoothness results as above follows from the non-smoothness of $V^{N}_{[Z]}$ for small values of $N$ like $2,3$ (the case $N=2$ has been partially treated in cohomological terms in [mclean2005]). The strongest evidence to 5.1 is the following theorem in [ho13, Theorem 19.1, part 2] which is proved by heavy computer calculations.
@@
In the context of 5.1, for r,rˇ, $\in\mathbb{N},\check{r}\in\mathbb{Z},\ 1\leq r,|\check{r}|\leq 10$, the infinitesimal Hodge locus $V_{[Z]}^{N},N\leq M$ is smooth for all $(n,M)=(6,14),(8,6),(10,4),(12,3)$.
For $n=4$, the Hodge locus $V_{[Z]}$ itself is smooth for trivial reasons. There is abundant examples of Hodge cycles for which we know neither to verify the Hodge conjecture (construct algebraic cycles) nor give evidences that they might be counterexamples to the Hodge conjecture, see [Deligne-HodgeConjecture] and [ho13, Chapter 19]. Finding Hodge cycles for hypersurfaces is extremely difficult, and the main examples in this case are due to T. Shioda for Fermat varieties [Shioda1979].